If you invest $3,500 for 3 years and 3 months, how much is it worth to the nearest dollar if it increases 7% a year?

N=No(1+-r)^t
N=3500(1.07)^?

Answer by checkley77(12844)

N=No(1+-r)^t
N=3500(1.07)^3.25 (3 MONTHS IS ONE QUARTER OF A YEAR).
N=3500(1.24594)
N=$4,360.79 OR $ 4,361 IS THE VALUE OF THE INVESTMENT @ 7% FOR 3 YEARS & 3 MONTHS.